### v1

# GOAL: A Generalist Combinatorial Optimization Agent Learner

Abstract
Machine Learning-based heuristics have recently shown impressive performance in solving a variety of hard combinatorial optimization problems (COPs). However they generally rely on a separate neural model, specialized and trained for each single problem. Any variation of a problem requires adjustment of its model and re-training from scratch. In this paper, we propose GOAL (for Generalist combinatorial Optimization Agent Learning), a generalist model capable of efficiently solving multiple COPs and which can be fine-tuned to solve new COPs. GOAL consists of a single backbone plus light-weight problem-specific adapters, mostly for input and output processing. The backbone is based on a new form of mixed-attention blocks which allows to handle problems defined on graphs with arbitrary combinations of node, edge and instance-level features. Additionally, problems which involve heterogeneous nodes or edges, such as in multi-partite graphs, are handled through a novel multi-type transformer architecture, where the attention blocks are duplicated to attend only the relevant combination of types while relying on the same shared parameters. We train GOAL on a set of routing, scheduling and classic graph problems and show that it is only slightly inferior to the specialized baselines while being the first multi-task model that solves a variety of COPs. Finally we showcase the strong transfer learning capacity of GOAL by fine-tuning or learning the adapters for new problems, with only few shots and little data.
1 Introduction
Combinatorial Optimization problems are extensively studied in the Operations Research community, because they are both challenging in terms of theoretical complexity and central in many applications such as transportation, logistics and finance. Traditional methods include exact solvers (typically based on integer programming [32]) and approximation algorithms [55] which provide optimality or approximation guarantees, respectively, but do not scale; and heuristic algorithms [9] which produce good-quality solutions faster but without guarantees. Effective heuristics heavily rely on expert knowledge and are generally problem-specific. A more recent and promising family of methods for CO is based on machine learning, either to learn the hyper-parameters of classic CO solvers, or to directly learn heuristics in an end-to-end fashion [6, 11]. In particular constructive neural heuristics have been successful for solving a variety of CO problems, including routing [5, 42, 30, 33], scheduling [58, 34] and graph problems [27, 45]. However, in most cases, a separate model is specialized and trained for each problem, and sometimes even for each distribution of instances within a problem, especially for varying instance sizes. While much effort has been devoted to improving the generalization to different instance distributions [41, 59, 49, 16], generalization to different problems is less studied. Indeed, although there are strong neural heuristics for many CO problems, any variation of these problems requires adjustment of their model and re-training from scratch. While it is true that some problem-specific model tailoring is always required, since the structure of problem instances and solutions are very diverse, many problems are related and it would be interesting to exploit their commonalities. In this paper, we seek to answer the following question: Can one model be pretrained in order to be easily specialized to solve a variety of CO problems?
The idea of pretraining a unified model on some pretext task so that it can be efficiently fine-tuned to several downstream tasks has been extremely successful in the language and vision domains [15, 10, 46]. In these domains, a self-supervised pretraining task exploits the intrinsic structure and regularities of images and text to learn generic representations. In contrast, in CO, while most problem instances can be modeled as graphs, there is no generic regularity to leverage by a self-supervised task. Graphs, unlike images or text, are a “poor” modality, with little redundancy or intrinsic semantics.
In this paper, we propose GOAL (for Generalist combinatorial Optimization Agent Learning), a generalist model capable of efficiently solving a variety of CO problems and which can be fine-tuned to solve new related problems. GOAL consists of a single backbone and light-weight problem-specific adapters for input and output processing. To handle the large discrepancy between CO problems, we propose two complementary strategies. First, the backbone is based on a new form of mixed-attention blocks which allows to handle problems defined on graphs with arbitrary combinations of node, edge and instance-level features. Additionally, problems which involve heterogeneous nodes or edges, which typically arise with multipartite graphs, are handled through a novel multi-type transformer architecture, where the backbone mixed-attention blocks are duplicated to attend each relevant combination of types while relying on the same shared parameters. GOAL is able to handle any CO problem whose instances can be represented as a graph and where feasible solutions can be constructed iteratively, as a sequence of node selections. We find that this pattern is generic enough to cover a wide range of problems. We train GOAL in a multi-task setting, by imitation of expert trajectories of eight classic CO problems, spanning multiple domains: Routing problems: the Asymmetric Traveling Salesman Problem (ATSP), the Capacitated Vehicle Routing Problem (CVRP), the CVRP with Time Windows (CVRPTW), the Orienteering Problem (OP); Scheduling problems: the Job Shop Scheduling Problem (JSSP) and the Unrelated Machine Scheduling Problem (UMSP); Packing problems: the Knapsack Problem (KP); and Graph problems: the Minimum Vertex Covering Problem (MVC). The expert trajectories are based on good-quality solutions, easily generated by traditional problem-specific solvers for instances of medium size (generally 100 nodes).
Experimentally we first show that GOAL performs very well on all the training tasks, making it the first model that solves such a variety of CO problems. Second, we illustrate GOAL’s strong zero (or few) shot generalization performance to various distributions shifts from the training tasks’ instances. Third, we fine-tune GOAL on four new problems and demonstrate that it achieves a good performance quickly using little data. We also discuss scenarios where fine-tuning is less effective, hinting at the required conditions for successful transfer; namely that the new task is close enough to at least one of the training tasks to share a similar instance structure, while its objective and constraints can be completely different. We further provide an ablation study to evaluate the effectiveness of our multi-type transformer architecture and mixed-attention blocks. Notably our architecture gives state-of-the-art results on 6 of our 8 training tasks when used in a single-task version of GOAL. In summary, our contributions are the following:
  • We introduce GOAL: a generic and flexible model for end-to-end combinatorial optimization. Two key ingredients of GOAL are the simple yet effective form of mixed-attention blocks which enables handling problems represented by graphs with any combination of node, edge and instance-level features and its multi-type transformer architecture which allows to seamlessly tackle problems with multipartite graph structures.
  • We demonstrate that it is possible to learn a single backbone model to effectively solve a variety of CO tasks, spanning routing, scheduling, packing and graph problems.
  • We showcase the potential of such a multi-task pretrained model to be efficiently adapted to handle new CO tasks and provide empirical insights on how closely related the new tasks should be to the training tasks to ensure successful transfer.
2 Related Works
End-to-end heuristic learning for CO
Our work falls into the category of end-to-end constructive heuristics for combinatorial optimization. Constructive neural heuristics have been successfully applied to a wide range of COPs, including many variants of routing problems [30, 57, 28, 31, 18, 52], scheduling problems [58, 34], as well as classic COPs such as the Knapsack Problem [33] and the Maximum Independent Set [27, 52]. Another successful category of neural CO methods are improvement heuristics, which focus on iteratively enhancing a given complete solution, typically by using a neural network to guide the selection of local search operators [13, 22, 40, 56]. The seminal work of [27] introduces graph embeddings which do work for a variety of CO problems but with limited performance. Current state-of-the-art models are tailored for certain problem structures (e.g. Euclidian routing problems, where nodes are represented by their coordinates plus some extra node or graph features such as demand or vehicle capacity), and successfully handled by transformer-based architectures [30]. For non-euclidian routing problems, [34] and [16] propose adapted architectures to deal with the distance matrix, but they are specialized for such problems and do not apply to those with node-only features. In GOAL, we propose a novel mixed-attention block which can flexibly deal with both node and edge features, incurring no overhead when one or the other is missing.
Generalization of neural heuristics
While the first seminal models for neural CO had a dramatic drop in performance when tested on out-of-training distribution instances [26, 41], several recent approaches show a greatly improved generalization performance, especially to larger instances [18, 35, 49, 43], to distribution shifts of other instance parameters [8], or both [41, 8, 59, 16]. Concurrently with our work, some papers have started to explore the cross-problem generalization for the family of routing problems. Specifically, [36] propose to pretrain the AM model [30] on the TSP and to fine-tune either the entire model or some light-weight modules on the Price Collecting TSP and OP. Additionally, [60] and [37] explore generalization to different variants of the CVRP. They build on the idea of zero-shot compositionality [48] by training a model on CVRPs with some combinations of constraints from a predefined set (e.g. time windows and duration limit constraints) and show that the model generalizes to some unseen combinations of the same constraints. Although these approaches tackle multiple problems, it is important to note that the structures of the considered Euclidian routing problems are very similar and indeed many of those problems were solved by the seminal Attention Model of [30] with little adaptations of the base model. Our paper goes further and considers multi-task training and transfer learning of a new model on a variety of COPs with different structure.
Multi-task training for sequential decision making
Multi-task (pre-)training has recently shown promising results in related fields. For example, [25] proposed a graph neural network processor capable of learning to execute a variety of algorithms, such as sorting and searching. The paper focuses only on polynomial problems and learns to imitate the execution of classical algorithms. In contrast, we address NP-hard problems and learn by imitation of the solution provided by (potentially exponential) oracle algorithms. Recent approaches in control rely on sequence modeling of expert trajectories from multiple tasks [50, 51]. Although they treat various tasks, with potentially diverse state and action spaces, these works still consider image observations and exploit this structure. [47] proposes GATO, a general-purpose agent obtained by training a large transformer-based model on a hundreds of tasks spanning vision, language and control. It demonstrates the potential of a single network (albeit with 1.2B parameters) to effectively tackle a wide variety of tasks.
3 Multi-Task Learning for Combinatorial Optimization
Refer to caption Refer to caption
Figure 1: Left: Architecture of GOAL for single-type problems; the green components are task-specific, while the others belong to the backbone. Right: Architecture of GOAL for multi-type problems (with here 2 types of nodes T1,T2); “self” and “cross” are identical multi-head mixed attention blocks (same parameters for all types and tasks); “FF” for the two types are also identical Feed Forward blocks; for the edges, all the type combinations are shown (here T11,T12,T21,T22) but each task can define its own. The same set of parameters is used for the two architectures.
3.1 Constructive CO problem solving
We assume that a CO problem instance is given by a pair of a finite, non-empty set $X$ of feasible solutions, and a real-valued objective function $f$ whose domain contains $X$. Solving the instance consists in finding an optimal solution, i.e. an element of $\arg\min_{x\in X}f(x)$. Constructive approaches assume that $f$ is defined not only over $X$ but over a set of partial solutions, which contains $X$ and is amenable to a step-wise construction procedure: each partial solution can be obtained by a sequence of construction steps. Thus, solving the problem instance can be formulated as a sequential decision procedure which iteratively selects the construction steps so as to reach not only a feasible but an optimal solution. That in turn can be modeled as a Markov Decision Problem (MDP) associated with the problem instance.
The action space of this MDP is naturally defined as the (discrete) set of construction steps, but there are many ways to choose its state space. For the very broad class of so called “tail-recursive” CO problems, [16] shows that an effective choice is the space of all possible instances of the problem. Indeed, the construction of a solution can be modeled as a trajectory in the space of problem instances, following the tail sub-problems [7] obtained after each construction step. All the classical standard NP-hard CO problems, such as vehicle routing problems, machine scheduling, various graph problems etc. are tail-recursive, sometimes modulo a slight reformulation (e.g. TSP is not per se tail-recursive, but path-TSP is). In the sequel, we deal exclusively with tail-recursive CO problems and their construction MDPs, called BQ-MDP in [16].
3.2 Multi-Task CO Problems
Now, given a class $\mathcal{T}$ of tail-recursive CO problems called tasks, let $\boldsymbol{\Omega}_{t}$ for each task $t{\in}\mathcal{T}$ denote the parametric space used to represent its instances. Consider the new (multi-task) CO problem defined as the union of the tasks in $\mathcal{T}$ (each being a set of instances). Its instances can be represented in the disjoint union $\biguplus_{t\in\mathcal{T}}\boldsymbol{\Omega}_{t}$ consisting of pairs $(t,\omega)$ of a task identifier $t{\in}\mathcal{T}$ and a point $\omega{\in}\boldsymbol{\Omega}_{t}$ in its corresponding parameter space. It is easy to show that the multi-task problem is tail-recursive whenever each of its components is. And its BQ-MDP is then simply the disjoint union of the BQ-MDPs of its components.
In this paper, we study alternatives for the multi-task policy space, beyond the trivial one where no parameter is shared between the policies of the different tasks (Cartesian product of the policy spaces of the individual tasks). The underlying assumption is that if the tasks are sufficiently similar, learning one task may elicit implicit skills useful to other tasks, which are then leveraged by parameter sharing. In the extreme, some tasks may require little training data and/or just a few shot fine-tuning to obtain acceptable performance, leveraging their similarity to more thoroughly trained tasks. Since each task has its own parametric space used to represent its instances, some part of any multi-task policy has to be task specific, and trained by processing at least a few instances of each task: true zero-shot generalization is not possible. The situation is close to multi-lingual Natural Language Processing, where the vocabulary of each language needs to be seen at least once at training. And we borrow the solution used in that context, namely adapters [4], to capture the task-specific parts of the policy.
3.3 Task Representation
To leverage operational similarity between tasks, we need them to have at least representational similarity. Thus, we assume that each instance of size $N$ of a task (CO problem) is defined by a set of $N$ abstract “nodes”, together with features which can be attached at node, edge (node pair) or instance level. Instance features can be replicated and appended to the features of each node, hence can be ignored without loss of generality. Since the construction steps (actions) are returned by the task-specific output adapters (Sec 3.4), there are no restrictions on their format. For simplicity and to cover the problems used in our experiments, we assume in the sequel that each construction step consists of the selection of a node, possibly with an option among a fixed number $K$ of options (dependent on the task: in most cases, $K{=}1$ and the action simply selects a node).
Hence, the GOAL model takes three inputs $t,x,\bar{x}$, where $t$ is a task identifier, and $(x,\bar{x}){\in}\boldsymbol{\Omega}_{t}$ is a parameter representing an instance of task $t$, where $x{\in}\mathbb{R}^{N{\times}F_{t}}$ (resp. $\bar{x}{\in}\mathbb{R}^{N{\times}N{\times}\bar{F}_{t}}$) represents the feature vectors of dimension $F_{t}$ of the $N$ nodes (resp. $\bar{F}_{t}$ of the $N{\times}N$ edges). While some tasks may have no edge feature ($\bar{F}_{t}{=}0$), we always assume that at least one node feature exists, which distinguishes the nodes by a randomly assigned value between 00 and $1$, acting as node identifier, so we always have $F_{t}{\geq}1$. Both $F_{t}$ and $\bar{F}_{t}$, as well as the meaning of the corresponding features, depend on task $t$. The output of the model is a matrix $y{\in}\mathbb{R}^{N{\times}K_{t}}$ representing the score of each candidate action, i.e. the pair of a node and an option for selection. As usual, it is passed to a softmax (over the whole matrix) to obtain a probability distribution. Wherever appropriate, task-dependent masking can be used to ensure that probabilities are null for the actions disallowed by a task.
For example, for the path-TSP task, the nodes are the customers to visit plus the origin and destination of the vehicle. There are two proper node features, each holding a binary value indicating whether the node is the origin or the destination, respectively (thus $F_{t}{=}3$ counting node identifier). And there is a single edge feature ($\bar{F}_{t}{=}1$) holding the edge cost (distance, time, energy, etc.). The action is a simple node selection ($K_{t}{=}1$), with origin and destination nodes being masked. In a BQ-MDP transition, the selected node becomes the new origin, and the previous origin is discarded ($N$ is decreased by $1$).
3.4 Architecture of the Model
The architecture of our model generalizes that of BQ-NCO [16], and is described in Fig. 1 left. It essentially consists of a sequence of layers whose parameters are shared by all the tasks (backbone) together with task-specific adapter modules at the input and output of the model [^1] Additional adapters inside layers can also be introduced, but this paper considers only input-output adapters.. The input-output adapter modules for task $t$ are specific to the dimensions $F_{t},\bar{F}_{t},K_{t}$ and the meaning of the corresponding features. No positional encoding is used (all relational information is assumed to be captured by $\bar{x}$). The layers themselves, on the other hand, are only aware of the shared embedding dimensions $D,\bar{D}$. They are self-attention transformer layers [54] where vanilla attention is replaced by multi-head mixed attention, similar in intent to [34] but with a different design.
Our mixed attention block, in the general case where $M$ nodes attend on $N$ nodes ($M{=}N$ in self-attention), is an instance of generalized attention [2]. It takes the following inputs: node representations as vectors for $N$ queries, $M$ keys, and $M$ values, with an optional $M{\times}N$ log-binary mask $\mathcal{M}$, as in vanilla attention, plus optional edge representations as $M{\times}N$ “edge” vectors. In each head $h$, each query, key, value vector is first projected, as in vanilla attention, into a common space of dimension $d$, yielding for each node $n{=}1{:}N$ and $m{=}1{:}M$, vectors:
$Q_{n}^{(h)}=Q_{n}\mathbf{W}_{Q}^{(h)},\hskip 14.22636ptK_{m}^{(h)}=K_{m}\mathbf{W}_{K}^{(h)},\hskip 14.22636ptV_{m}^{(h)}=V_{m}\mathbf{W}_{V}^{(h)}.$
The edge vectors are also projected into the same space of dimension $d$, but twice so: once as a key and once as a query, resulting in vectors:
$Q^{\prime(h)}_{mn}=E_{mn}\mathbf{W}^{\prime(h)}_{Q},\hskip 14.22636ptK^{\prime(h)}_{mn}=E_{mn}\mathbf{W}^{\prime(h)}_{K}.$
Vanilla attention then builds an $M{\times}N$ score matrix $S_{mn}^{(h)}{=}Q_{n}^{(h)}K_{m}^{(h)\top}$, from which the attention coefficients are computed (using softmax and involving the mask). We simply augment the factors of the scalar product defining the score matrix with the edge projections, and keep the rest unchanged, yielding new representations $r$ of the $N$ nodes:
$\begin{split}r=\sum_{h}\textrm{softmax}_{\textrm{col}}(S^{(h)}{+}\mathcal{M})^{\top}V^{(h)}\mathbf{W}_{O}^{(h)\top}\hskip 2.84544pt\\ \text{ with }\hskip 2.84544ptS^{(h)}_{mn}{=}(Q_{n}^{(h)}{+}Q^{\prime(h)}_{mn})(K_{m}^{(h)}{+}K^{\prime(h)}_{mn})^{\top}.\end{split}$ (1)
In GOAL, all the node inputs are of dimension $D$, the edge inputs of dimension $\bar{D}$, and they are all projected into a space of dimension $d$ equal to $D$ divided by the number of heads. Furthermore, no masking is used in the attention blocks.
One well-known limitation of this model is the quadratic complexity bottleneck of attention matrices, which are of size $N^{2}$. In practice though, for many tasks, a simple, task-specific, heuristic can deal with that problem. Essentially, instead of presenting all the $N$ nodes of the input instance to the model, the heuristic selects a priori a subset of the $n$ most promising ones. Thus, the attention matrices never exceed size $n^{2}$, even when $N^{2}$ would be unmanageable. The performance impact is often negligible, and even sometimes improving (maybe due to the dilution of attention coefficient when $N$ is large). For example, with TSP, only the origin, destination and 100-nearest neighbors to the origin are presented.
3.5 Multi-Type Transformer Architecture
Many tasks (CO problems) involve different types of nodes. For example, scheduling problems often involve “operation” nodes (the things to schedule) and “machine” nodes (which constrain feasible schedules). For GOAL, one solution is to simply treat them all indiscriminately in the backbone. The adapters, being problem specific, are of course free to treat them differently according to their types. This is sometimes unsatisfactory though, since it forces the adapter to provide uniform embeddings for all the edges, though they are often heterogeneous in the input: in the scheduling example, edge features may be precedence constraints between operations, or processing times between machines and operations, and no information at all between machines. Instead of packing all these different types of relations into a single big tensor, with padding values wherever a feature is not defined for a given edge, we propose a multi-type architecture where the adapters output separate embeddings per type (for nodes) or compatible pairs of types (for edges). Each layer of the model, instead of being a monolithic self-attention block operating on all the nodes, becomes a task-specific combination of both cross- and self-attention blocks, as described by (1), operating on different type-compatible subsets of nodes and edges. This is illustrated in Fig. 1 right. Therefore, in our scheduling example, the model will contain separate self-attention blocks, for operation nodes and machine nodes respectively, and cross-attention blocks, for operation-machine edges and vice-versa. Note that even when edges are not oriented, it is useful to have two cross-attention blocks with transposed matrices as edge embeddings: the projections in (1) are then used to interpret differently the two directions of the edges in each attention block. Finally, observe that while the architecture configuration is now task-specific, the model parameters remain the same for all the possible configurations. This is precisely what enables us to train a unique model to solve various single and multi-type problems.
4 Experiments
Our experiments seek to answer the following questions: Does the proposed GOAL approach enable learning a single backbone to effectively solve a variety of CO problems? (Sec 4.1); Can we use it as a pretrained model and efficiently fine-tune it to solve new unseen CO problems? (Sec 4.2); and finally, to what extent do the introduced mixed-attention blocks and multi-type architecture contribute to the results? (Sec 4.3). We start by describing the multi-task training of GOAL.
Training tasks.
We train GOAL on eight classic and varied CO problems. For each problem, we generate a training dataset of 1 million random instances and solve them using traditional solvers. Specifically, we consider the following problems and oracle solvers: the Asymmetric Traveling Salesman Problem (ATSP) with LKH [19], the Capacitated Vehicle Routing Problem (CVRP) with LKH [19], the Capacitated Vehicle Routing Problem with Time Windows (CVRPTW) with HGS [19] the Orienteering Problem (OP) with A4OP [29], the Job Shop Scheduling Problem (JSSP) with ORTools [44], the Unrelated Machine Scheduling Problem (UMSP) with HiGHS [24], the Knapsack Problem (KP) with ORTools [44] and the Minimum Vertex Covering Problem (MVC) with FastWVC [1]. The training instances are of size 100 nodes, or 10$\times$10 for JSSP and 10$\times$100 for UMSP (resulting in both cases in 100 operation nodes and 10 machine nodes). For the sake of space, the description of the problems and associated data generation processes is provided in Appendix A.
Training process.
GOAL is trained by alternating batches from each of the training tasks (round robin), using a single-class cross entropy loss for ATSP, CVRP, CVRPTW, OP, JSSP and UMSP and a multi-class cross entropy loss for KP and MVC. Similarly to [16], we create a batch by sampling suffix subproblems for each instance. We use AdamW optimizer [38], with an initial learning rate of 0.0005 and a decay rate every 10th epochs. We trained the model on 8 Nvidia A100 GPU servers, with batch size of 128 for 7 days, allowing the completion of approximately 500 epochs. The best model was selected based on the average performance on a validation set of 512 instances per problem. The model hyperparameters are described in Appendix C.
Baselines.
We consider the following constructive neural models as baselines: AM [30], MDAM [57], POMO [33], Sym-NCO [28], BQ-NCO [16], MVMoE [60], MatNet [34], S2V-DQN [27] and COMPASS [12]. These models are all specialized and trained in single-task mode. While all these models are usually combined with a search procedure, such as beam search [16], simulation-guided beam search  [14], instance augmentation [33], active search [23], for a fair comparison of the underlying neural models, and for sake of space, we only report here the results with a greedy application of each model.
4.1 Performance on the Training Tasks
[…]
ATSP100 CVRP100 CVRPTW100 OP100
opt. gap time opt. gap time opt. gap time opt. gap time
Oracle solver 0.00% 29s 0.00% 12m 0.00% 10m 0.00% 1.1m
AM greedy - 7.11% 1s 5.35% 1s
MDAM greedy - 4.84% 1s - 2.88% 16s
POMO no aug - 1.21% 1s - -
Sym-NCO greedy - 3.33% 1s - 2.03% 2s
BQ-NCO greedy 1.27% 2s 2.79% 2s - 0.22% 2s
MVMoE/4E - 1.65% 1s 4.90% 1s -
MatNet greedy 0.93% 1s - - -
GOAL single-task greedy 0.29% 5s 2.10% 5s 2.54% 5s 0.08% 3s
GOAL multi-task greedy 0.45% 5s 2.85% 5s 3.21% 5s 0.51% 3s
KP100 MVC100 UMSP100x10 JSSP10x10
opt. gap time opt. gap time opt. gap time opt. gap time
Oracle solver 0.00% 1s 0.00% 2m 0.00% 2m 0.00% 12m
POMO no aug 0.19% - - -
BQ-NCO greedy 0.10% - - -
S2V-DQN - 0.97% 2s - -
COMPASS - - - 4.70% 23m
GOAL single-task greedy 0.10% 3s 0.31% 3s 1.82% 7s 2.65% 6m
GOAL multi-task greedy 0.12% 3s 0.34% 3s 2.28% 7s 3.23% 6m
Table 1: Performance on 128 test instances from the training tasks. Lower is better.
Generalization to new instance distributions.
Since GOAL is trained with supervised learning on instances of a certain size, it is critical that its performance generalizes to larger instances. Table 2 presents the generalization to instances up to 10 times bigger than the training ones, for a subset of the training tasks. We see an excellent generalization performance on ATSP and CVRP. For the JSSP, the generalization is not as strong but still very good on most distributions. Note that scaling to much larger instances is not a focus of this paper. Recent approaches which do not rely on quadratic attention blocks such as [39] would be more appropriate, although highly specialized to some problems. We believe that the more expensive mixed-attention blocks are crucial for the model’s capacity to handle diverse tasks. We also tested finetuning GOAL to other distribution shifts, namely to the Euclidian distance for TSP and asymmetric distances for CVRP (while GOAL was trained on ATSP and Euclidian CVRP). Results are presented in Appendix B and demonstrate a strong performance.
Train. distribution Generalization
ATSP100 ATSP200 ATSP500 ATSP1000
MatNet greedy 1.57% 1s 97.21% 12s - -
BQ-NCO greedy 1.63% 1s 1.55% 9s 4.29% 26s 8.09% 1m
GOAL multi-task greedy 0.45% 5s 1.54% 30s 1.48% 3m 1.89% 1h
CVRP 100 CVRP200 CVRP500 CVRP1000
AM greedy 7.11% 1s 10.16% 1s 17.2% 1s 72.95% 2.1s
MDAM greedy 4.84% 1s 7.54% 6s 12.94% 24s 32.25% 1m
POMO no aug 1.21% 1s 5.94% 6s 31.36% 32s 284.58% 1m
Sym-NCO greedy 3.33% 1s 8.42% 3s 37.58% 54s 481.53% 9m
BQ-NCO greedy 2.79% 1s 2.81% 9s 3.64% 55s 5.88% 2m
GOAL multi-task greedy 2.85% 5s 2.80% 12s 3.63% 48s 6.63% 3m
JSSP10x10 JSSP15x15 JSSP20x15 JSSP20x20
COMPASS 4.70% 23m 8.00% 38m 10.40% 1h -
GOAL multi-task greedy 2.65% 6m 5.90% 18m 17.40% 32m 10.80% 59m
Table 2: Generalization performance on larger instances. Lower is better.
4.2 Fine-Tuning to New Tasks
In this section, we want to evaluate the transfer capability of the trained GOAL backbone to unseen tasks. We consider four new CO problems: the Traveling Repairman Problem (TRP), the Open Vehicle Routing Problem (OVRP), the Split Delivery Vehicle Routing Problem (SDVRP) and the Maximum Independent Set (MIS) problem (detailed description in Appendix A). For each problem, our first attempt was to define new task-specific input and output adapters and train them either with or without back-propagation through the backbone. It turned out that in that case, the fine-tuning process is very unstable and has a strong dependency on the parameter initialization (i.e. the random seed). Next, we initialized the input adapters with the ones of the closest training problems: TSP for TRP, CVRP for OVRP and SDVRP and MVC for MIS. The output adapters were still trained from scratch. In this case the fine-tuning is very effective. In Fig 2, we see that the finetuned model reaches very good optimality gaps (at most $5\%$) much faster than a single-task GOAL model trained from scratch. Our few-shot fine-tuning process relies on only 64 instances and lasts less than 1 minute. We use the same loss and optimizer as for the training but with a smaller initial learning of 0.0001. These results assume that the whole model is fine-tuned, including the backbone. When the latter is frozen, we observe a slight degradation of the results with no significant speedup. From these experiments, we hypothesize that the input adapters are the difficult part of the transfer. To test this hypothesis, we considered one of the training tasks, discarded its input and/or output adapters and tried to re-learn them from scratch, with the trained backbone. This led to similar observations: re-learning the output adapter works, while re-learning the input adapter is very unstable. We conclude that transfer in GOAL is successful for problems that share the same instance structure as one of the training problems and thus can benefit from a pre-trained input adapter. Note that the new problems can have arbitrary objective and constraints (as is the case with the four tested new tasks). Appendix D further discusses input adapters and their impact on the latent representations.
Figure 2: Fine tuning GOAL to new tasks.
4.3 Ablation Study
Figure 3: Comparison of the model performance when using different variants of mixed-attention blocks for solving the Asymmetric Traveling Salesman problem
Figure 4: Comparison between single and multi-type architectures for the Job Shop Scheduling Problem
Mixed-attention blocks.
To evaluate the effectiveness of our mixed-attention blocks, we replace them by alternative modules which also aim at mixing the edge information with the attention scores, namely MatNet [34] and G2G [20]. More precisely we focus on the ATSP and train three versions of GOAL with the three different blocks. For MatNet, we keep the original hyper-parameters used for solving the ATSP. Figure 4 presents the models performance (optimality gap on a test set of 128 instances) at different points of the training. We see that using our mixed-attention blocks leads to a significant gain, on the performance at convergence (MatNet) or on the speed of convergence (G2G). Note that these ablations are performed on a smaller training dataset than in the previous experiments (100k instances), hence the wider opt gap.
Multi-type transformer.
The proposed multi-typed transformer architecture is meant to deal with tasks that are naturally represented by multipartite graphs with heterogeneous nodes or edges, typically scheduling problems where jobs and machine nodes play different roles. To evaluate its effect on the performance, we compare it to that obtained with the more direct alternative which ignores node heterogeneity: we simply concatenate the features of the different types of nodes (or edges) and use dummy values whenever a feature is undefined, so that the original single-type model can be used. We implemented this alternative for the Job Shop Scheduling Problem and present in Fig 4 a comparison between the training performance of the single-type and multi-type architectures. Clearly the multi-type architecture converges faster and to a better performance: after 200 epochs of training on 100K instances, it reaches a test optimality gap of 4.6% while the single-type model achieves only 8.9%.
5 Discussion and Conclusion
In this paper, we propose a generic neural CO model (GOAL) and demonstrate the feasibility of learning a shared backbone model to solve a variety CO problems, spanning routing, scheduling, packing and classic graph problems. We have shown the benefits of GOAL as (i) a single-task specialized model, providing state-of-the-art results on six CO problems; (ii) a multi-task model, able to solve several problems with the same backbone; and (iii) a pretrained model, to be efficiently fine-tuned for new problems. To the best of our knowledge, this is the first neural approach that handles such a variety of problems with a single backbone model. We believe this is a key step in the direction of a foundation model for CO, where the diversity of problem representations (as opposed to the uniformity of textual or visual data) is still a major hurdle.
Our experiments have exclusively been conducted with Imitation Learning, which assumes the availability of good-quality solutions for at least a certain size of instances. While this is not an intrinsic limitation of the GOAL model – the underlying MDPs being defined independently of the training process – we believe that Imitation Learning is an effective training paradigm for learning multi-task models. This is indeed demonstrated by previous works which relied on sequence-modeling training tasks for control [51, 47]. Leveraging solutions to classical NP-hard problems, provided by traditional solvers, which are based on decades of OR research, is particularly advantageous for pretraining a model. Such a model can later be adapted to more exotic problems with minimal supervision data, or even better, with Reinforcement Learning. Indeed using Reinforcement Learning to fine-tune GOAL to new problems, for which high-quality solvers may not be available, constitutes a promising direction for future work.
Appendix A Description of the considered CO problems
In this Section, we describe the CO problems that are solved by GOAL as well as the instance generation process. We organize the problems into three categories: Routing, Scheduling and Packing and Graph Problems. We also provide an overview Table of the problems and used solvers for each category.
A.1 Routing Problems
TSP TRP OP
Node features - - Price
Edge features Distance between nodes
Instance features - - Maximum tour length
Objective $min$(tour length) $min$(sum of waiting times) $max$(collected prices)
Oracle Concorde(LKH) LKH EA4OP
Action Select a node
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
CVRP/SDCVRP CVRPTW OCVRP
Node features Demands Demands Time window starts Time window ends Service times Demands
Edge features Distance between nodes
Instance features Vehicle capacity
Objective Minimize total tour length
Oracle Concorde/HGS HGS LKH
Action Select a next node, subroute is done
[Uncaptioned image] [Uncaptioned image]
Table 3: Overview of the routing problems solved by GOAL
A.1.1 Traveling Salesman Problem
The Traveling Salesman Problem (TSP) is a classic combinatorial optimization problem. The objective of the TSP is to find the shortest possible route that a salesman can take to visit a given set of cities exactly once and return to its starting point. In the path-TSP variant, the origin and destination of the salesman are distinct nodes (but may have the same coordinates, so that the TSP is a special case of the path-TSP). A TSP instance can be represented either by node coordinates or by a distance matrix. The majority of NCO models address TSP in the coordinate setting, yet the matrix setting holds greater significance for practical applications, as it applies beyond the Euclidean context, where the matrix can represent time, energy or cost, rather than distance, and need not satisfy the symmetry property of a distance matrix.
Data generation
For the Euclidian case, we generate datasets as described in [30]: points are uniformly sampled from the unit square, and the corresponding distance matrix $A$ is computed. For non-Euclidean problems (ATSP) we generate data as described in [34], where distance matrices are generated randomly, but post-processed to satisfy the triangular inequality.
A.1.2 Capacitated Vehicle Routing Problem
The Capacitated Vehicle Routing Problem (CVRP) is another standard routing problem. The goal of CVRP is to optimize the delivery routes of a fleet of vehicles to a set of customers, while considering each vehicle’s capacity. The objective is to minimize the overall distance traveled by the vehicles, while respecting the capacity constraints and ensuring that all the customer demands are met.
Data generation.
The matrix is sampled in the same way as for TSP, while node demands are randomly generated integers from the interval [1, 10]. Following [30], the vehicle capacity depends on the problem size: capacity of 50 for a problem size of 100, 80 for 200, 100 for 500, and capacity of 250 for problems with 1000 nodes.
A.1.3 CVRP with Time Windows
The CVRP with Time Windows (CVRPTW) is an extension of the CVRP. In addition to demand constraints, each customer also has a specified time window during which they can be visited, along with the necessary service time required for serving that customer. The objective of the CVRPTW is to find routes for the vehicles that satisfy the capacity limits, deliver goods within the specified time windows, and minimize the total travel and waiting time or the number of vehicles used. In our model, we consider the standard version of the problem where total traveling and serving time should be minimized.
Data generation.
For the time windows and service times, we use a similar procedure to [17], where the travel times between nodes are sampled in the range [0, $\sqrt{2}\pm 0.1$]. We ensure the feasibility of the resulting instances by sampling time windows using a more refined procedure. Windows start time values are sampled from the range (0.01, 0.25). It’s important to note that windows start time do not directly impact problem feasibility, as vehicles can wait at the node until the service starts. On the other hand, the choice of the windows end time has an impact on feasibility. An instance is unfeasible if the window end time of a customer is less than the travel time from the depot plus the service time. To avoid this, windows end time are sampled from the interval [max(window start time, travel time from depot) + 2 $\times$ service time, max(window start time, travel time from depot) + 4 $\times$ service time].
A.1.4 Split Delivery VRP
The Split Delivery VRP (SDVRP) is similar to the CVRP with the difference that a customer can be visited by several vehicles, i.e. its demand is split between multiple vehicles.
Data generation.
We use the same data as for the CVRP (Sec A.1.2).
A.1.5 Open Vehicle Routing Problem
This is a variant of CVRP where an unbounded fleet of vehicles is available at the depot, each serving a set of customers, and they do not need to return to the depot at the end of their service.
Data generation.
We use the same data as for the CVRP (Sec A.1.2).
A.1.6 Orienteering Problem
The Orienteering Problem (OP) is a combinatorial optimization problem that involves finding the optimal route to maximize the collected prizes while visiting a set of locations and satisfying a maximal distance constraint.
Data generation.
Again we follow previous work [30]: node locations are sampled in the same way as for the TSP and CVRP and prizes are randomly sampled integers from the interval [1, 10]. The maximal distance constraint is set to 4.
A.1.7 Traveling Repairman Problem
In the travelling repairman problem (TRP), as in TSP, a set of locations is given as well as the time needed to travel between any pair of locations. Each location has a task to be performed by the repairman. The objective of the problem is to find the route that minimizes the sum of waiting times for all tasks.
Data generation.
We use the same data as for the TSP (Sec A.1.1)
A.2 Packing and Graph Problems
KP MVC MIS
Node features Price and weight -
Edge features - Adjacency matrix
Instance features Knapsack capacity -
Objective $max$(collected prices) $max$(vertex coverage) $min$(independent set)
Oracle ORTools FastWVC
Action Select a node
Table 4: Overview of the packing and graph problems solved by GOAL
A.2.1 Knapsack Problem
The Knapsack Problem (KP) is a well-known combinatorial optimization problem that involves selecting a subset of items from a given set. Each item had a weight and a value, and the goal is to maximize the cumulated value of selected items, but without exceeding a given cumulated weight threshold, known as the knapsack’s (weight) capacity.
Data generation.
Following the same procedure as in [33], node weights and values are randomly sampled from the unit interval and knapsack capacity is set to 25 for all problem sizes.
A.2.2 Minimum (Weighted) Vertex Cover Problem
The goal of the Minimum Vertex Cover Problem (MVC) is to find the smallest set of vertices in a graph, such that each edge is incident to at least one of these vertices. In the weighted version of this problem (MWVC), each vertex has an associated weight, and the goal is to select vertices with the minimum cumulated sum of weights while satisfying the vertex cover condition.
Data generation.
For this problem, we follow a similar procedure as in [27]. We generate random graphs by using Erdős–Rényi model with the critical edge probability $p$ sampled from [0.05, 0.15].
A.2.3 Maximum Independent Set
In the Maximum Independent Set (MIS), we are given an undirected graph $G=(V,E)$ and the goal is to find a subset of nodes $S\subset V$ of maximal size $|S|$ such that each pair of nodes in $S$ is not connected by an edge.
Data generation.
We use the same data generation process as for the Minimum (Weighted) Vertex Cover Problem.
A.3 Scheduling Problems
UMSP JSSP
Node features Machine: Availability time Task: Execution time, Availability time Machine: Availability time
Edge features Job-machine: Execution times Task-task: precedence constraints, $ij=1$ if tasks $i$ and $j$ belong to the same job Task-machine: $ij=1$ if task $i$ must be executed on machine $j$
Objective Minimize a makespan
Oracle HiGHS ORTools
Action Select a machine for a random task Select a task
[Uncaptioned image] [Uncaptioned image]
Table 5: Overview of the scheduling problems solved by GOAL
A.3.1 Job Shop Scheduling Problem
The Job Shop Scheduling Problem (JSSP) is a classic scheduling problem. We are given $N$ jobs and $M$ machines. Each job consists of a sequence of operations that need to be processed in a specific order (implying precedence constraints). Each operation has a given processing time. Each machine can only process one operation at a time and when an operation is started it cannot be interrupted (no preemption). A solution is a schedule, defined by the start time of each operation such that the makespan, i.e. the maximum completion time of the operations, is minimized.
Data generation.
We follow the standard procedure for generating JSSP benchmark datasets: execution times are randomly sampled from $U(1,10)$.
A.3.2 Unrelated Machine Scheduling Problem
In the Unrelated Machine Scheduling problem (UMSP), the goal is to schedule $N$ jobs on $M$ machines such that the makespan is minimized. Each job $i$ has a processing time $p_{ij}$ on each machine $j$, which is arbitrary (hence the term unrelated). There are no precedence constraints and the schedules are non-preemptive. The solution to the problem is essentially a partition of the jobs on the machines. [21] shows that this problem is NP-hard.
Data generation.
For UMSP, the processing times are randomly generated from $U(1,100)$.
Appendix B Finetuning of GOAL on Various Distribution Shifts
We consider the following distribution shifts: TSP with Euclidian distances and CVRP with asymmetric distances (while GOAL was trained on ATSP and Euclidian CVRP). Fig 5 compares finetuning GOAL on these new distributions, versus training a new single-task model from scratch. We observe that finetuning converges faster and to a better optimality gap.
Figure 5: Tuning GOAL to various distribution shifts
Appendix C GOAL Hyperparameters
The GOAL backbone consists of 9 layers with an embedding size of 128, a feed-forward dimension of 512, and ReZero normalization [3]. The mixed-attention blocks have 8 heads.
Appendix D Analysis of the latent representations
To investigate the interaction between the different CO problems, we compare the latent representations of the different tasks and their evolution through the layers of the trained GOAL model. Since these are high-dimensional vectors, we use the standard t-SNE technique [53] to visualize the representations in the 2D space. More precisely, in Fig 6 we plot the t-SNE projection (with perplexity $150$) of the node embeddings after the first, fourth, and last layer of the model, for each of the training tasks (top) and with the addition of the transfer tasks (bottom). For each task, we plot 20 full instances of size 100 (110 for scheduling problems). We observe that the representations seem clustered per problem in the different layers. We also see clear overlaps between similar problems (CVRP-SDCVRP-OVRP in green shades and TSP-TRP in blue shades) in layers 1 and 4 while they become separated in the last layer. These visualizations indicate that the problem-specific input adapters seem to map the instances to distinct areas of the embedding space. This might explain the strong dependency between the input adapters and the backbone which makes learning a new input adapter for a new task often challenging.
Figure 6: Visualization of the evolution of the node latent representations through the GOAL model for the training tasks (top) and with the addition of the transfer tasks (bottom)